OpenStudy (imnotgoodatmath):

Match Term Definition 11^0 A)-49 (-49)^1 B) 1/49 0^1 C) 0 7-2 D) 1

OpenStudy (anonymous):

11^0=1 (-49)^1=-49 0^1=0 7^-2=1/49